A pharmacist has two solutions of alcohol, a 15% solution and a 28% solution. How many milliliters of each solution should be mixed to gather to make a 100ml solution which is 20% alcohol
Answer by mananth(16946)

Alcohol
percent ---------------- quantity
alcohol I 15.00% ---------------- x ml
Alcohol 2 28.00% ------ 100 - x ml
Mixture 20.00% ---------------- 100
Total 100 ml
15.00% x + 28.00% ( 100 - x ) = 20.00% * 100
15 x + 28 ( 100 - x ) = 2000
15 x + 2800 - 28 x = 2000
15 x - 28 x = 2000 - -2800
-13 x = -800
/ -13
x = 61.54 ml 15.00% alcohol I
38.46 ml 28.00% Alcohol 2

m.ananth@hotmail.ca